OpenStudy (anonymous):

SOLVING EQUATIONS WITH VARIABLES ON EACH SIDE -7(2b-4)=5(-2b+6) I keep getting; -35 = 35, so does that mean the problem has no solution. I just wanna double check so I'm not wrong on my homework!

OpenStudy (shamil98):

-7(2b-4)=5(-2b+6) -14b +28 = -10b + 30 4b = 2 b= 1/2

OpenStudy (shamil98):

Oh oops, -14b + 28 = -10b + 30 -4b = 2 b= -1/2

OpenStudy (anonymous):

How is that because the first equals 21 and the second equals 25.

OpenStudy (shamil98):

What do you mean the first one equals 21?..

OpenStudy (anonymous):

I just typed it into my calculator. -7(2(1,2)-4 = 21

OpenStudy (shamil98):

-7(2b-4) =5(-2b+6) Distributive property of multiplication. -14b+28=-10b + 30 -14b = -10b + 2 -4b = 2 b = - 1/2

OpenStudy (shamil98):

-1/2 = -0.5

OpenStudy (shamil98):

-7(2-(0.5)-4) = 5 (-2-(0.5) +6) -7(-1-4) = 5(1+6) -7(-5) = 5(7) 35 = 35

OpenStudy (shamil98):

All you need to know for your answer is that B = -1/2 , okay? inputting the value of b just makes sure b works in this problem got it?